# Domain-Adaptive Pre-training of Self-Supervised Foundation Models for Medical Image Classification in Gastrointestinal Endoscopy

### v1

@@ Abstract
Video capsule endoscopy has transformed gastrointestinal endoscopy (GIE) diagnostics by offering a non-invasive method for capturing detailed images of the gastrointestinal tract, enabling early disease detection. However, its potential is limited by the sheer volume of images generated during the imaging procedure, which can take anywhere from 6-8 hours and often produce up to 1 million images, necessitating automated analysis. Additionally, the variability of these images, combined with the need for expert annotations and the scarcity of large, high-quality labeled datasets, constrains the effectiveness of current medical image analysis models. To address this, we introduce a novel large gastrointestinal endoscopy dataset, called EndoExtend24, created by merging and re-stratifying the train/test splits of ten existing public and private datasets, ensuring no overlap of patient data across splits. EndoExtend24 includes over 226,000 labeled images, as well as dynamic class mappings, which allow unified training across datasets with differing labeling granularity, supporting up to 123 distinct pathological findings. Further, we propose to leverage domain adaptive pre-training of foundation models in computer vision trained with self-supervision on generic image data, to adapt them to the task of GIE medical diagnosis. Specifically, the EVA-02 model, which is based on the vision transformer architecture and was trained on ImageNet-22k with masked image modeling (using EVA-CLIP as a MIM teacher), is pre-trained on the novel EndoExtend24 dataset to achieve domain adaptation, and finally trained on the Capsule Endoscopy 2024 Challenge dataset. Experimental results show promising results on the challenge validation set, with an AUC Macro score of 0.993 and a balanced accuracy of 89.3%.
@@ 4 Evaluation
The performance of our final eva02-base model was evaluated on the CE24 validation set, consisting of 10,658 images. Table 2 presents a comparative analysis of our model’s performance against baseline models SEER and VGG16, both after pre-training and on the final downstream task. The eva02-base model demonstrates superior performance across all metrics in both stages. Notably, in the downstream task, it achieves a balanced accuracy of 0.893, a macro AUC of 0.993, a weighted F1 score of 0.960, and a macro mAP of 0.931. Figure 3 provides a visual representation of the model’s classification performance, as indicated by receiver operating characteristic (ROC) curves.
@@
The left panel displays individual ROC curves for each of the 10 distinct pathological findings, with AUC values ranging from 0.94 to 0.99. This indicates strong discriminative ability across all gastrointestinal findings. The right panel shows the micro- and macro-average ROC curves, both achieving an AUC of 0.99 and 0.97 respectively, which further corroborates the model’s robust overall performance in multi-class classification. The collective evidence indicates that the eva02-base model effectively utilizes adaptive pre-training and demonstrates resilience in addressing the distinctive challenges inherent to GIE multi-class classification, as exemplified by the CE24 dataset.
@@ 5 Conclusion
In this work, we presented a novel approach for domain-adaptive pre-training of self-supervised foundation models in the context of GIE diagnostics. By introducing the EndoExtend24 dataset, a comprehensive collection of over 226,000 labeled images across multiple GIE modalities, we addressed key challenges such as dataset scarcity, terminology inconsistency, and data leakage. Our experiments demonstrate the effectiveness of pre-training the EVA-02 model on EndoExtend24 and fine-tuning it on the Capsule Endoscopy 2024 Challenge dataset, achieving impressive results with a Macro AUC score of 99.33 and a balanced accuracy of 89.32 on the challenge validation set. These results highlight the potential of leveraging large-scale, domain-adaptive self-supervised pre-training to enhance the performance of foundation models for specialized downstream tasks such as video capsule endoscopy. Our future work will explore further improvements to dataset standardization and model interpretability to support broader adoption of AI in clinical settings and foster the explainability of model results for enhanced clinical insights.

### v2

@@ Abstract
Video capsule endoscopy has transformed gastrointestinal endoscopy (GIE) diagnostics by offering a non-invasive method for capturing detailed images of the gastrointestinal tract, enabling early disease detection. However, its potential is limited by the sheer volume of images generated during the imaging procedure, which can take anywhere from 6-8 hours and often produce up to 1 million images, necessitating automated analysis. Additionally, the variability of these images, combined with the need for expert annotations and the scarcity of large, high-quality labeled datasets, constrains the effectiveness of current medical image analysis models. To address this, we introduce a novel large gastrointestinal endoscopy dataset, called EndoExtend24, created by merging and re-stratifying the train/test splits of ten existing public and private datasets, ensuring no overlap of patient data across splits. EndoExtend24 includes over 226,000 labeled images, as well as dynamic class mappings, which allow unified training across datasets with differing labeling granularity, supporting up to 123 distinct pathological findings. Further, we propose to leverage domain adaptive pre-training of foundation models in computer vision trained with self-supervision on generic image data, to adapt them to the task of GIE medical diagnosis. Specifically, the EVA-02 model, which is based on the vision transformer architecture and was trained on ImageNet-22k with masked image modeling (using EVA-CLIP as a MIM teacher), is pre-trained on the novel EndoExtend24 dataset to achieve domain adaptation, and finally trained on the Capsule Endoscopy 2024 Challenge dataset. Experimental results demonstrate strong performance with an F1 score of 0.88, an improvement of about 39% over the baseline model’s F1 score of 0.49. This substantial improvement can be attributed to the fact that the baseline models are relatively primitive compared to the state-of-the-art models employed in our approach. Additionally, the model achieved a macro AUC score of 0.993 and a balanced accuracy of 89.3%.
@@ 4 Evaluation
The performance of our final eva02-base model was evaluated on the CE24 validation set, consisting of 10,658 images. Table 2 presents a comparative analysis of our model’s performance against baseline models SEER and VGG16, both after pre-training and on the final downstream task. The eva02-base model demonstrates superior performance across all metrics in both stages.
@@
Notably, in the downstream task, it achieves a balanced accuracy of 0.893, a macro AUC of 0.993, a weighted F1 score of 0.960, and a macro mAP of 0.931. Figure 3 provides a visual representation of the model’s classification performance, as indicated by receiver operating characteristic (ROC) curves. The left panel displays individual ROC curves for each of the 10 distinct pathological findings, with AUC values ranging from 0.94 to 0.99.
@@
Figure 4: Confusion matrix of our fine-tuned eva02-base on the CE validation set.
This indicates strong discriminative ability across all gastrointestinal findings. The right panel shows the micro- and macro-average ROC curves, both achieving an AUC of 0.99 and 0.97 respectively, which further corroborates the model’s robust overall performance in multi-class classification. The collective evidence indicates that the eva02-base model effectively utilizes adaptive pre-training and demonstrates resilience in addressing the distinctive challenges inherent to GIE multi-class classification, as exemplified by the CE24 dataset. Figure 4 shows the confusion matrix for the multi-class classification of pathological findings, with the highest accuracy observed for normal mucosa, with 8001 true positives, and worms, with 44 true positives. Misclassifications are most frequent between similar classes like erosion and normal mucosa, while rare conditions like worms show minimal confusion with other classes.
@@ 5 Conclusion
In this work, we presented a novel approach for domain-adaptive pre-training of self-supervised foundation models in the context of GIE diagnostics. By introducing the EndoExtend24 dataset, a comprehensive collection of over 226,000 labeled images across multiple GIE modalities, we addressed key challenges such as dataset scarcity, terminology inconsistency, and data leakage. Our experiments demonstrate the effectiveness of pre-training the EVA-02 model on EndoExtend24 and fine-tuning it on the Capsule Endoscopy 2024 Challenge dataset, achieving impressive results with an F1 score of 0.88, a Macro AUC score of 99.33 and a balanced accuracy of 89.32 on the challenge validation set. These results highlight the potential of leveraging large-scale, domain-adaptive self-supervised pre-training to enhance the performance of foundation models for specialized downstream tasks such as video capsule endoscopy. Our future work will explore further improvements to dataset standardization and model interpretability to support broader adoption of AI in clinical settings and foster the explainability of model results for enhanced clinical insights.